Solve using the substitution method. If the system has no solution or an infinite number of solutions, state this.
-x + 6y = 54
3x + y = 9
Solve using the elimination method. If the system has no solution or an infinite number of solutions, state this.
-2x - 8y = 22
-6x - y = 20

Answer by mananth(12270)

-x+6y=54 .............1
y=9-3x ---------2
y=9-3x
Plug the value of y in (1)
-1x+6(9-3x)=54
-1x+54 -18x=54
-1x-18x =54-54
-19x=0
x= 0
Plug the value of x in (1)
-x+6y=54
-1*0+6y=54
6y=54
6y=54
/6
y=9
------------------------------
-2x-8y= 22 .............1
-6x-y= 20 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -8
-2x-8y=22
48x++8y=-160
Add the two equations
46x=-138
/46
x=-3
plug value of x in (1)
-2x-8y= 22
6-8y=22
-8y=22-6
-8y=16
y=-2